# FENIX: Enabling In-Network DNN Inference with FPGA-Enhanced Programmable Switches

### v1

@@ 7 EVALUATION
Our evaluation addresses several key questions: (i) What classification accuracy can FENIX achieve on real traffic datasets, and how does it compare to other methods? (ii) What computational resources are required by FENIX, including both the programmable switching ASICs during the feature extraction phase and the FPGA resources? (iii) How is the performance of FENIX affected under different scales of concurrent flows and throughput? (iv) As an asynchronous hybrid system, does FENIX introduce noticeable latency overhead when using the FPGA for inference?
@@ 7.1 Methodology
(a) FENIX-CNN-FLOW: We implemented a flow-level CNN with 3 convolutional layers (64, 128, 256 filters) and 2 fully connected layers (512, 256 neurons). Flow accuracy is calculated based on majority voting of packet classifications within each flow.
(b) FENIX-RNN-FLOW: We developed a flow-level RNN with a single custom RNN cell (128 units) and a dense output layer. The model processes packet length and IPD features through embeddings and classifies flows using the final hidden state.
@@
Tasks. We use the following tasks to evaluate FENIX, as summarized in Table 1: (i) Encrypted Traffic Classification on VPN: This task aims to classify network traffic that has been encrypted by VPNs. We use the ISCXVPN2016 dataset [19]. (ii) Malware Identification: This task distinguishes between traffic generated by benign applications and various types of malware. We use the USTC-TFC2016 dataset [47].
Metrics. We use flow-level macro-F1 (the average of F1-scores across different classes) as the accuracy metric for FlowLens, while for other schemes we report packet-level macro-F1. Additionally, we provide a breakdown of Precision and Recall for each class to enable more detailed performance analysis.
@@ 7.2 Classification Accuracy on Real Datasets
7.3 Hardware Resource Utilization
Table 3 compares the hardware resource overhead of several representative P4 systems on programmable switches. The results show that FENIX achieves relatively low resource usage across multiple dimensions. For instance, its SRAM consumption (12.9%) is much lower than FlowLens (34.2%), indicating more efficient memory management. In terms of TCAM usage, FENIX also stands out, with only 4.4% overhead compared to NetBeacon’s 18.8%, highlighting that FENIX avoids the trade-off of reducing one resource at the cost of sharply increasing another. Beyond memory, FENIX maintains efficient use of bus bandwidth and pipeline stages, remaining below or comparable to most other systems. This balanced and predictable resource profile is enabled by the decoupled DataEngine design, which keeps FENIX ’s overhead stable across different tasks and workloads.
Figure 10: Scaling test of FENIX
@@ 7.4 Flow Count and Throughput Scalability
We conduct stress tests on FENIX under high-concurrency and high-throughput scenarios. Since all the original network trace files were collected in low-bandwidth networks (e.g., tens of Mbps), we generate high-throughput trace files by concurrently packaging a large number of flows (ensuring each flow has a unique identifier) and reassigning new timestamps to these packets  [^2] We write the original timestamp into the packet header to handle changes caused by flow acceleration.. We then replay these pcap files on our testbed to generate actual network traffic.
Figure 10 presents the results of the scalability test, where we progressively increase the flow concurrency until the physical bandwidth of the traffic generator’s NIC is fully saturated. The results demonstrate that FENIX can easily handle this level of scale, as its macro-F1 scores remain nearly identical to those in Table 2.
To evaluate FENIX at even larger scales, we build a simulator to emulate the entire workflow of FENIX. By replicating the experimental settings of Table 2 and Figure 10, we validate the accuracy of the simulator, finding that the simulation results are almost the same as those from our testbed. Subsequently, we use the simulator to explore significantly larger scales, progressively increasing the flow concurrency to tens of thousands of flows per second and the aggregate throughput to the Tbps level. As shown in Figure 10, the macro-F1 score of FENIX experiences only a minor decrease, with a reduction of about 13.2% at the largest scale.
Figure 11: Average Latency Compare of FENIX and FlowLens.
@@
Most importantly, FENIX reduces inference latency by nearly three orders of magnitude compared to FlowLens. While FlowLens requires over $1,000\mu$s for each inference, FENIX completes inference in just $1.2\mu$s on average. This improvement comes from in-network acceleration on FPGA, which offloads model execution from the host CPU and enables efficient, microsecond-scale inference.
@@ 8 Discussion
Hardware Adaptability Beyond Tofino. FENIX is prototyped using a Tofino 2 switch and Xilinx ZU19EG FPGA, but its architecture is fundamentally portable. The discontinuation of Intel’s Tofino line [23] does not threaten the broader applicability of FENIX, as the programmable switch ecosystem remains rich, including Juniper Trio [52], Broadcom Trident [20], and RMT-based research designs such as dRMT [13] and Menshen [46]. Industry trends—exemplified by Broadcom Trident 5’s NetGNT, which integrates neural traffic analysis into switching silicon—indicate that intelligent data planes are becoming mainstream, though open programming interfaces for such capabilities remain limited. The hardware and software co-design of FENIX anticipates industry trends and provides a platform that can be easily adapted to new generations of programmable switches.
@@
Automation and Future Directions. Recent frameworks such as Homunculus [45] highlight the growing need for automated ML pipelines and model deployment in data plane environments. While FENIX currently requires expert-driven integration when migrating to new tasks or hardware, its modular Model Engine design is well-suited for future automation. Extending FENIX with high-level, task-agnostic abstractions and automated toolchains for FPGA and switch targets would further lower deployment barriers and accelerate its adoption across diverse operational scenarios.
@@ 9 Related Work
ML for Network Traffic Analysis. Machine learning has become a core tool for network traffic analysis, powering solutions for malicious traffic detection [30, 6, 18, 34], website fingerprinting [36, 14, 38], and fine-grained traffic classification [4, 39]. The field is rapidly advancing toward more sophisticated models and tasks, including traffic prediction and encrypted flow analysis [28, 54, 56]. However, deploying these models at line rate in the data plane remains challenging due to stringent hardware constraints—motivating innovations in model compression, quantization, and hardware-aware co-design.
Data Plane Device Roles and Limitations. Modern network infrastructures rely on programmable switches, SmartNICs, and FPGAs—each with distinct strengths. SmartNICs efficiently offload compute from servers for functions such as rate limiting, traffic analysis, and protocol acceleration [35, 40, 41, 9]. Programmable switches provide unparalleled throughput and enable in-network tasks such as DDoS defense and real-time classification [5, 29, 50, 27, 51, 57, 26, 2, 24, 49, 12, 3]. FPGAs serve as both prototyping platforms and as flexible data plane accelerators, often complementing the rigid resource budgets of ASIC-based switches [42, 44, 16, 37, 53].
Positioning of FENIX. Prior work has either focused on deploying quantized or highly compressed ML models in programmable switches—often sacrificing model complexity and accuracy—or has relied on executing high-precision models at the network edge, typically with limited throughput and incomplete visibility. By contrast, FENIX explores a new point in the design space: it leverages the tight integration of FPGA and programmable switch hardware at central network nodes to support efficient, high-throughput DNN inference in the data plane, achieving a practical trade-off between inference accuracy, adaptability, and deployment feasibility. This hybrid approach demonstrates that advanced neural inference capabilities can be meaningfully integrated into switch-centric architectures without requiring full line-rate DNN execution, offering a pragmatic path toward more intelligent and flexible network data planes, and informing future ASIC and automation developments.
@@ 10 CONCLUSION
Appendix A Problem Formulation
Variable Description
$V$ Token generation rate
$Q$ Global packet rate
$N$ Number of flows
$T_{i}$ Duration of backlogged packets
$C_{i}$ Number of backlogged packets
$Q_{i}$ Current flow packet rate
$P_{i}$ Probability of generating feature vector
Table 5: Variables of token generation probability model.
As described in the main text, our token allocation model is based on two fundamental criteria:
Criterion 1: If all flows have the same rate, to achieve fair allocation, each flow’s transmission period should be $\frac{N}{V}$.
Criterion 2: In practical scenarios where flow rates differ, tokens should be allocated proportionally to flow rates, resulting in a transmission period of $\frac{Q}{Q_{i}V}$ for each flow, where $Q_{i}=\frac{C_{i}}{T_{i}}$ represents the current flow’s packet rate.
Based on these criteria, we develop a probability model that determines when a flow should send feature information to the FPGA. The model ensures that flows of different rates have appropriate opportunities to transmit, while maintaining overall system stability.
For a flow with packet rate $Q_{i}=\frac{C_{i}}{T_{i}}$, we establish a time-dependent linear probability model that varies according to the relationship between $\frac{N}{V}$ and $\frac{Q}{Q_{i}V}$. The mathematical expression for this probability model is as follows:
$P_{i}(T_{i},C_{i})=\begin{cases}\frac{C_{i}(VT_{i}-N)}{QT_{i}-NC_{i}},&\text{if }T_{i}\in\left[\frac{N}{V},\frac{Q}{Q_{i}V}\right],\left(\frac{N}{V}<\frac{Q}{Q_{i}V}\right)\\ \frac{T_{i}(VC_{i}-Q)}{NC_{i}-QT_{i}},&\text{if }T_{i}\in\left[\frac{Q}{Q_{i}V},\frac{N}{V}\right],\left(\frac{N}{V}>\frac{Q}{Q_{i}V}\right)\\ 1,&\text{if }QT_{i}=NC_{i}\text{ and }T_{i}\geq\frac{N}{V}\\ 0,&\text{if }QT_{i}=NC_{i}\text{ and }T_{i}<\frac{N}{V}\end{cases}$ (3)
To verify this model theoretically, we analyze the expected transmission periods. For any flow $i$, the probability of sending feature information over time can be represented as a cumulative distribution function:
$P(t)=\frac{t-t_{s}}{t_{e}-t_{s}}\quad(t\in[t_{s},t_{e}])$ (4)
The expected value of transmission time is:
[…]
where $f(t)$ is the probability density function, $f(t)=P^{\prime}(t)=\frac{1}{t_{e}-t_{s}}$.
This gives us $E(t)=\frac{t_{s}+t_{e}}{2}$.
For both cases ($\frac{N}{V}<\frac{Q}{Q_{i}V}$ and $\frac{N}{V}>\frac{Q}{Q_{i}V}$), the expected period for flow $i$ is:
$E_{i}=\frac{Q_{i}N+Q}{2Q_{i}V}$ (6)
Taking the average of expected periods across all flows, we have:
$E$ $=\frac{\sum_{i=1}^{N}Q_{i}\cdot E_{i}}{Q}$ (7)
$=\frac{\sum_{i=1}^{N}Q_{i}\cdot\frac{Q_{i}N+Q}{2Q_{i}V}}{Q}$ (8)
$=\frac{N\sum_{i=1}^{N}Q_{i}+Q\sum_{i=1}^{N}1}{2QV}$ (9)
$=\frac{NQ+QN}{2QV}$ (10)
$=\frac{N}{V}$ (11)
This result confirms that our model achieves fair token resource allocation on average, balancing the needs of both high-rate and low-rate flows while preventing FPGA overload, which aligns with our design objectives as outlined in the main text.

### v3

@@ 7 EVALUATION
Our evaluation addresses several key questions: (i) What classification accuracy can FENIX achieve on real traffic datasets, and how does it compare to other methods? (ii) What computational resources are required by FENIX, including both programmable switching ASICs during feature extraction and FPGA resources? (iii) How is FENIX performance affected under different scales of concurrent flows and throughput? (iv) As an asynchronous hybrid system, does FENIX introduce noticeable latency overhead when using FPGA for inference?
@@ 7.1 Methodology
(a) FENIX-CNN-FLOW: We implemented a flow-level CNN with 3 convolutional layers and 2 fully connected layers. Flow accuracy is calculated based on majority voting of packet results within each flow.
(b) FENIX-RNN-FLOW: We developed a flow-level RNN with a single custom RNN cell and a dense output layer. The model processes packet length and IPD features through embeddings and classifies flows using the final hidden state.
@@
Tasks. We use the following tasks to evaluate FENIX, as summarized in Table 1. (i) Encrypted Traffic Classification on VPN: This task aims to classify network traffic that has been encrypted by VPNs. We use the ISCXVPN2016 dataset [25]. (ii) Malware Identification: This task distinguishes between traffic generated by benign applications and various types of malware. We use the USTC-TFC2016 dataset [57]. Note that model training is performed offline, and the FPGA is used exclusively for inference tasks.
Metrics. We use flow-level macro-F1 as the accuracy metric for FlowLens, while for other schemes we report packet-level macro-F1. We also provide a breakdown of Precision and Recall for each class to enable detailed performance analysis.
@@ 7.2 Classification Accuracy on Real Datasets
7.3 Hardware Resource Utilization
Table 3 compares the hardware resource overhead of several representative P4 systems on programmable switches. The results show that FENIX achieves relatively low resource usage across multiple dimensions. For instance, its SRAM consumption (12.9%) is much lower than FlowLens (34.2%), indicating that the introduction of FPGA effectively reduces the resource overhead of P4 systems. In terms of TCAM usage, FENIX also stands out, with only 4.4% overhead compared to NetBeacon’s 18.8%, highlighting that FENIX avoids the trade-off of reducing one resource at the cost of sharply increasing another. Beyond memory, FENIX maintains efficient use of bus bandwidth and pipeline stages, remaining below or comparable to most other systems. This balanced and predictable resource profile is enabled by the decoupled DataEngine design, which keeps FENIX ’s overhead stable across different tasks and workloads.
@@ 7.4 Flow Count and Throughput Scalability
We conduct stress tests on FENIX under high-concurrency and high-throughput scenarios. Since original network trace files were collected in low-bandwidth networks, we generate high-throughput trace files by concurrently packaging flows with unique identifiers and reassigning timestamps. Figure 10 shows results where we progressively increase flow concurrency until the traffic generator’s NIC bandwidth is saturated. FENIX easily handles this scale, with macro-F1 scores nearly identical to Table 2.
To evaluate larger scales, we build a simulator emulating FENIX’s workflow. We configure maximum concurrent flow count based on switch chip registers and set inference latency according to measured results. After validating simulator accuracy against our testbed, we explore significantly larger scales with tens of thousands of flows per second and Tbps-level throughput. As shown in Figure 10, FENIX’s macro-F1 score experiences only a minor decrease, with about a 13.2% reduction at the largest scale.
@@
Figure 11: Latency Compare of FENIX and FlowLens.
@@ 8 Discussion
Hardware Adaptability Beyond Tofino. FENIX is prototyped using a Tofino 2 switch and Xilinx ZU19EG FPGA, but its architecture is fundamentally portable. The discontinuation of Intel’s Tofino line [32] does not threaten the broader applicability of FENIX, as the programmable switch ecosystem remains rich, including Juniper Trio [65], Broadcom Trident [29], runtime programmable switch designs [21, 61, 22], and RMT-based research designs such as dRMT [15] and Menshen [56]. For example, Broadcom Trident 5’s NetGNT integrates neural traffic analysis directly into switching silicon. However, publicly available information on such commercial products lacks detailed specifications of AI accelerator capabilities and programming interfaces, making it difficult for the research community to leverage these platforms for academic research and innovation. The hardware and software co-design of FENIX anticipates industry trends and provides a platform that can be easily adapted to new generations of programmable switches.
@@
Automation and Future Directions. Recent frameworks such as Homunculus [55] highlight the growing need for automated ML pipelines and model deployment in data plane environments [63]. While FENIX currently requires expert-driven integration when migrating to new tasks or hardware, its modular Model Engine design is well-suited for future automation. However, the current FPGA implementation has inherent limitations regarding flexibility. Layer parameters and feature dimensions are fixed during synthesis, meaning that any DNN architecture updates or changes typically require complete hardware re-synthesis, which limits system adaptability. Our current design primarily supports common CNN and RNN models. Dynamic model updates and partial reconfiguration mechanisms can further enhance system flexibility, and these represent future research directions. Extending FENIX with high-level, task-agnostic abstractions, automated toolchains for FPGA and switch targets, and runtime dynamic model switching capabilities would further lower deployment barriers and accelerate its adoption across diverse operational scenarios.
@@ 9 Related Work
ML for Network Traffic Analysis. Machine learning has become a core tool for network traffic analysis, powering solutions for malicious traffic detection [40, 6, 24, 44], website fingerprinting [46, 17, 48], and fine-grained traffic classification [4, 49]. The field is rapidly advancing toward sophisticated models and tasks, including traffic prediction and encrypted flow analysis [38, 68, 70]. However, deploying models at line rate in data plane remains challenging due to hardware constraints, motivating innovations in model compression, quantization, and hardware-aware co-design.
Data Plane Device Roles and Limitations. Modern network infrastructures rely on programmable switches, SmartNICs, and FPGAs—each with distinct strengths. SmartNICs efficiently offload compute from servers for functions such as rate limiting, traffic analysis, and protocol acceleration [45, 50, 51, 11, 58]. Programmable switches provide unparalleled throughput and enable in-network tasks such as DDoS defense and real-time classification [5, 39, 62, 36, 64, 71, 67, 35, 2, 33, 60, 14, 3]. FPGAs serve as both prototyping platforms and as flexible data plane accelerators, often complementing the rigid resource budgets of ASIC-based switches [27, 19, 28, 9, 52, 37, 54, 20, 47, 66].
Comparison and Positioning. Several commercial products integrate hardware accelerators (FPGAs, DPUs, or AI accelerators) with programmable switch ASICs, such as APS Networks’ APS6120Q [10] (Intel Tofino 3.2 Tbps ASIC with dual Intel Stratix 10 MX FPGAs) for broadband gateways and network security, and Cisco’s smart switch series [16] for enterprise networking. However, these products do not publicly disclose accelerator-switch integration details, deployable model scales, or resource allocation strategies for network intelligence tasks. FENIX differs by providing a complete co-design methodology specifically for high-throughput DNN inference in the data plane, addressing the throughput gap challenge between heterogeneous chips and conducting real-world evaluations. Prior work either deploys quantized models in switches (sacrificing accuracy) or executes high-precision models at network edges (introducing latency). FENIX leverages tight FPGA-switch integration at central nodes to achieve practical trade-offs between inference accuracy and speed, demonstrating that advanced neural inference can be effectively integrated into switch-centric architectures. This work provides design insights that could inform both commercial deployments and future ASIC developments for intelligent network data planes. We envision these commercial products could serve as potential deployment platforms for FENIX, further advancing intelligent network technologies.
@@ 10 CONCLUSION
Acknowledgements
We thank the anonymous reviewers and our shepherd, Eric Keller, for their insightful comments and suggestions. This work is supported by the National Science Foundation for Distinguished Young Scholars of China under Grant No. 62425201; the Science Fund for Creative Research Groups of the National Natural Science Foundation of China under Grant No. 62221003; the National Natural Science Foundation of China under Grant Nos. 62472240, 62394322, U22B2031, 62202473, and 62572473; the Taishan Scholar Foundation of Shandong Province under Grant No. tstp20250724; and the Beijing National Research Center for Information Science and Technology under Grant No. BNR2025RC01010. Ke Xu (xuke@tsinghua.edu.cn) and Su Yao (yaosu@tsinghua.edu.cn) are the corresponding authors.